OpenStudy (anonymous):

2x-10/9 find the graph and its inverse

OpenStudy (anonymous):

To find the inverse substitute x for y and y for x and then solve for y. Ex. y=2x-1 x=2y-1 1/2x=y-1 1/2x+1=y

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

got it shadly?

OpenStudy (anonymous):

@Shadly